OpenStudy (anonymous):

Write 5 log x + 4 log (x – 2) as a single logarithm.

OpenStudy (anonymous):

Use these : \[n\log a=\log a^n\\ \log a+\log b=\log(ab) \]

OpenStudy (anonymous):

A. 20 log x(x – 2) B. log x5(x – 2)4 C. log x(x – 2) D. none of these

OpenStudy (anonymous):

\[5\ln x+4\ln \left( x-2 \right)=\ln x ^{5}+\ln \left( x-2 \right)^{4}=\ln x ^{5}\left( x-2 \right)^{4}\]